OpenStudy (anonymous):

For questions 5-7, find the solution to the system of equations by using either graphing or substitution. 5. y=6-x and y=x-2 A. (2,4) B. (-4,2) C. (4,2) D No solutions

OpenStudy (anonymous):

\[6-x=x-2\iff 6=2x-2\iff 8=2x\iff x=4\]

OpenStudy (anonymous):

? Uhm what?

OpenStudy (anonymous):

your job it to solve for \(x\) you have \[6-x=x-2\] do you know how to solve it?

OpenStudy (anonymous):

No.. That's why I was asking for help.

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

Omg never mind. I think I figured it out! It's C. 4,2 right? Thanks